OpenStudy (anonymous):

. Write the equation of a line in point-slope form that has a slope of -4 and passes through the point (-7,6) explanation please

OpenStudy (texaschic101):

point slope form : y - y1 = m(x - x1) your slope (-4) is in the m position (-7,6)...x1 = -7 and y1 = 6 now just sub that into the formula y - 6 = -4(x - (-7) = y - 6 = -4(x + 7) <== your answer any questions ?

OpenStudy (anonymous):

thank you so much @texaschic101

OpenStudy (texaschic101):

no problem..do you have any questions ?

OpenStudy (anonymous):

no. but do can you help me with some more please @texaschic101

OpenStudy (texaschic101):

sure...what ya got

OpenStudy (anonymous):

Write the equation of a line that is perpendicular to and passes through (6, 3)

OpenStudy (texaschic101):

perpendicular to what ?

OpenStudy (texaschic101):

is there supposed to be an equation in there ? Or are you asking what is perpendicular to the last equation ?

OpenStudy (texaschic101):

hello ??

OpenStudy (anonymous):

no there is no equation but the answer is y=1/2x+6. can you explain @texaschic101